The sound level measured in a room by a person watching a movie varies from 50 dB to 100 dB during a loud part. Approximately how many times louder is the latter sound?

In a student laboratory exercise, the wavelength of a 30,000 Hz wave is measured to be 1.085 cm. Find the air temperature.
Answer #1

1,-50dB /2-100dB 1 = 12-11 = 100-50 = 50dB / 50 10 10 so the number of times the N 32 2) λ=1.085 × 10-2 m f =3x10Hz v=/1-3×1